OpenStudy (anonymous):

. Determine the slope of the line passing through the points (-7,7), (3,-5)

OpenStudy (anonymous):

FORMULA IS [Y-Y1]=m[X-X1] PUT THE VALUES; [7-(-5)]=m[-7-3] [7+5] =m[-10] [13] =m[-10] -13/10=m